# COAP: Memory-Efficient Training with Correlation-Aware Gradient Projection

### v1

@@ 2 Related Works
Memory-efficient Training Techniques are widely employed to reduce the overall memory footprint in large-scale model training, including activation checkpointing [4], 8-bit optimizers [8], and mixed-precision training [34], etc. The popular approach, ZeRO [38], optimizes memory usage in multi-GPU settings. ZeRO contains three stages, progressively sharding optimizer states, gradients, and model weights across devices, synchronizing only when necessary to minimize memory usage. Our method specifically targets memory reduction for optimizer states and can be effectively combined with these techniques, offering additional memory savings and further improving training efficiency.
@@ 3.3 Proposed Method
where $\mathrm{CosSim}(\cdot,\cdot)$ and $\mathrm{MSE}(\cdot,\cdot)$ return the cosine similarity and mean squared error, respectively. $\hat{\bm{G}}_{t}\in\mathbb{R}^{m\times n}=\bm{G}_{t}\bm{P}_{t-1}\bm{P}_{t-1}^{\top}$ and $\hat{\bm{M}}_{t-1}\in\mathbb{R}^{m\times n}=\bm{M}_{t-1}^{\rm proj}\bm{P}_{t-1}^{\top}$ are the full-rank estimates of gradient and first-order moment projected back from the low-rank subspace, respectively. Notably, the reconstruction term is introduced to minimize the reconstruction error for gradients incurred by projection – achieving the similar goal that SVD essentially aims for, and direction term encourages the consistency of optimization direction after restoring from low-rank subspace. To solve Eqn. 6 as a non-convex optimization problem, we use stochastic gradient descent (SGD) to iteratively update $\bm{P}_{t}$ (see details in Appendix).
@@
where $t=\lambda\times T_{\rm u}$ and $\Delta\bm{W}_{i}^{\rm proj}$ represent the weight update within the low-rank subspace defined by $\bm{P}_{i}$, with gradients over each $T_{\rm u}$ steps sharing the same $\bm{P}_{i}$. Notably, since $T_{\rm u}$ and $\lambda$ are hyper-parameters that control the update interval of $\bm{P}_{t}$, the corresponding ablation study on their effects is discussed in Section 5.
@@ 4 Experiments
In our experiments, we use various measures, and $\uparrow$” indicates that a higher value is better for this metric, while $\downarrow$” indicates that a lower value is better for this metric.
@@ 1 Detailed Proposed Method
$\min\limits_{\bm{P}_{t}}\underbrace{\mathrm{MSE}(\hat{\bm{G}}_{t},\bm{G}_{t})}_{\text{reconstruction term}}\underbrace{(1-\mathrm{CosSim}(\hat{\bm{M}}_{t-1},\bm{G}_{t}))}_{\text{direction term}},$ (1)
where $\mathrm{CosSim}(\cdot,\cdot)$ and $\mathrm{MSE}(\cdot,\cdot)$ return the cosine similarity and mean squared error, respectively. To simplify notation, the notation without the subscript $t$ represents a general form of the optimization problem, i.e.,
$\min\limits_{\bm{P}}\underbrace{\mathrm{MSE}(\hat{\bm{G}},\bm{G})}_{\text{reconstruction term}}\underbrace{(1-\mathrm{CosSim}(\hat{\bm{M}},\bm{G}))}_{\text{direction term}},$ (2)
@@
$\bm{P}:=$ $\bm{P}-\eta(\frac{\partial\mathrm{MSE}(\hat{\bm{G}},\bm{G})}{\partial\bm{P}}(1-\mathrm{CosSim}(\hat{\bm{M}},\bm{G})+$ (3)
$\quad\frac{\partial\mathrm{CosSim}(\hat{\bm{M}},\bm{G})}{\partial\bm{P}}\mathrm{MSE}(\hat{\bm{G}},\bm{G})).$
@@
$\frac{\partial\mathrm{MSE}(\hat{\bm{G}},\bm{G})}{\partial\bm{P}}$ $=\frac{\partial}{\partial\bm{P}}\left(\frac{1}{mn}\mathrm{tr}((\hat{\bm{G}}-\bm{G})^{\top}(\hat{\bm{G}}-\bm{G}))\right)$ (4)
@@ 2 Experimental Results
2.1 Pre-training DDPM
@@ 2.2 Qualitative Comparisons
We present qualitative results, showcasing images generated by models trained with COAP and other optimizers. These results provide a visual comparison that complements the quantitative analysis in the main paper. Comparisons for DDPM (CIFAR-10), DDPM (CelebA-HQ), LDM, SiT-XL/2, and ControlNet-XL are detailed in Tables 2345, and 6, respectively.

### v2

@@ 2 Related Works
Memory-efficient Training Techniques are widely employed to reduce the overall memory footprint in large-scale model training, including activation checkpointing [6], 8-bit optimizers [10], and mixed-precision training [36], zeroth-order optimization [2, 61, 4, 64], etc. The popular approach, ZeRO [40], optimizes memory usage in multi-GPU settings. ZeRO contains three stages, progressively sharding optimizer states, gradients, and model weights across devices, synchronizing only when necessary to minimize memory usage. Our method specifically targets memory reduction for optimizer states and can be effectively combined with these techniques, offering additional memory savings and further improving training efficiency.
@@ 3.3 Proposed Method
where $\mathrm{CosSim}(\cdot,\cdot)$ and $\mathrm{MSE}(\cdot,\cdot)$ return the cosine similarity and mean squared error, respectively. $\hat{\bm{G}}_{t}\in\mathbb{R}^{m\times n}=\bm{G}_{t}\bm{P}_{t-1}\bm{P}_{t-1}^{\top}$ and $\hat{\bm{M}}_{t-1}\in\mathbb{R}^{m\times n}=\bm{M}_{t-1}^{\rm proj}\bm{P}_{t-1}^{\top}$ are the full-rank estimates of gradient and first-order moment projected back from the low-rank subspace, respectively. Notably, the reconstruction term is introduced to minimize the reconstruction error for gradients incurred by projection – achieving the similar goal that SVD essentially aims for, and direction term encourages the consistency of optimization direction after restoring from low-rank subspace. To solve Eqn. 6 as a non-convex optimization problem, we use stochastic gradient descent (SGD) to iteratively update $\bm{P}_{t}$ (see theoretical analysis and implementation details in Appendix).
@@
where $t=\lambda\times T_{\rm u}$ and $\Delta\bm{W}_{i}^{\rm proj}$ represent the weight update within the low-rank subspace defined by $\bm{P}_{i}$, with gradients over each $T_{\rm u}$ steps sharing the same $\bm{P}_{i}$. Notably, since $T_{\rm u}$ and $\lambda$ are hyper-parameters that control the update interval of $\bm{P}_{t}$, the corresponding ablation study on their effects is discussed in Section 5. Additionally, we provide details on the CNN-specific handling and the Adafactor-based algorithm in Appendix.
@@ 4 Experiments
In our experiments, we use various measures, and $\uparrow$” indicates that a higher value is better for this metric, while $\downarrow$” indicates that a lower value is better for this metric. We provide the detailed hyper-parameter settings for all experiments in Section 1.3 of Appendix.
@@ 6 Conclusion
7 Acknowledgments
Special thanks are extended to Prof. Yuqian Zhang for providing the theoretical analysis that strengthened this research.
@@ 1 Detailed Proposed Method
$\min\limits_{\bm{P}_{t}}\underbrace{\mathrm{\rm MSE}(\hat{\bm{G}}_{t},\bm{G}_{t})}_{\text{reconstruction term}}\underbrace{(1-\mathrm{CosSim}(\hat{\bm{M}}_{t-1},\bm{G}_{t}))}_{\text{direction term}},$ (1)
where $\mathrm{CosSim}(\cdot,\cdot)$ and $\mathrm{\rm MSE}(\cdot,\cdot)$ return the cosine similarity and mean squared error, respectively. To simplify notation, the notation without the subscript $t$ represents a general form of the optimization problem, i.e.,
$\min\limits_{\bm{P}}\underbrace{\mathrm{\rm MSE}(\hat{\bm{G}},\bm{G})}_{\text{reconstruction term}}\underbrace{(1-\mathrm{CosSim}(\hat{\bm{M}},\bm{G}))}_{\text{direction term}},$ (2)
@@
$\bm{P}:=$ $\bm{P}-\eta(\frac{\partial\mathrm{\rm MSE}(\hat{\bm{G}},\bm{G})}{\partial\bm{P}}(1-\mathrm{CosSim}(\hat{\bm{M}},\bm{G})+$ (3)
$\quad\frac{\partial\mathrm{CosSim}(\hat{\bm{M}},\bm{G})}{\partial\bm{P}}\mathrm{\rm MSE}(\hat{\bm{G}},\bm{G})).$
@@
$\frac{\partial\mathrm{\rm MSE}(\hat{\bm{G}},\bm{G})}{\partial\bm{P}}$ $=\frac{\partial}{\partial\bm{P}}\left(\frac{1}{mn}\mathrm{tr}((\hat{\bm{G}}-\bm{G})^{\top}(\hat{\bm{G}}-\bm{G}))\right)$ (4)
@@
1.2 Theory Analysis.
COAP proposes to approximately keep the same low-rank subspace across iterations. Intuitively, this requires the low-rank subspace to remain stable across iterations, which can be bounded in terms of the learning rate and Lipchitz constant. Theoretical justification composed of the following steps:
  • Bounded update will lead to bounded change of the subspace.
  • Equation 2 finds a subspace not much worse than the true subspace.
We aim to demonstrate that this procedure guarantees an approximation to the underlying true subspace. Let $\bm{P}_{t}^{\natural}$ be the true subspace. Let $\bm{P}_{t}^{\rm MSE}$ and $\bm{P}_{t}^{\rm SIM}$ be the solutions of the MSE loss and CosSim loss respectively, then
$\bm{P}_{t}^{\rm MSE}$ $=\bm{P}_{t-1}$ (8)
$\bm{P}_{t}^{\rm SIM}$ $=\text{argmax}\left\langle\bm{G}_{t}\bm{P}_{t},\bm{M}_{t-1}^{proj}\right\rangle,$
where $\bm{M}_{t-1}^{\rm proj}=\beta\bm{M}_{t-2}^{\rm proj}+(1-\beta)\bm{P}_{t}^{T}\bm{G}_{t}=\sum_{i=0}^{t-1}\beta^{i}(1-\beta)\bm{P}_{t-i}^{T}\bm{G}_{t-i}$, $\text{argmax}\left\langle\cdot,\cdot\right\rangle$ maximizes an inner product between two entities. Therefore,
$\bm{P}_{t}^{\rm SIM}$ $=\sum_{i=0}^{t-1}\beta^{i}(1-\beta)\text{argmax}\left\langle\bm{P}_{t}^{T}\bm{G}_{t},\bm{P}_{t-i}^{T}\bm{G}_{t-i}\right\rangle$ (9)
$=\sum_{i=0}^{t-1}\beta^{i}(1-\beta)\bm{P}_{t-i}$
Theorem 1.1.
Assume that the gradient is Lipchitz with respect to the weight with Lipchitz constant $L$. Assume $\kappa_{1}\doteq\frac{\sigma_{1}}{\sigma_{r+1}}>1$ and $\kappa_{r}\doteq\frac{\sigma_{r}}{\sigma_{r+1}}>1$ to be the ratio between the $1$-th and $r$-th to the $r+1$-th singular value of the gradient $\bm{G}$ respectively. Then the proposed procedure in Equation 2 recovers a subspace well approximating the true subspace.
$\left\|\bm{P}_{t}-\bm{P}_{t}^{\natural}\right\|_{F}$ (10)
$~{}~{}\leq c(t\mod(\lambda\times T_{u}))\frac{r\kappa_{1}+\min(m,n)-r}{\kappa_{r}-1}L\eta.$
With learning rate $\eta$, Lipchitz property of the gradient gives
$\left\|\bm{G}_{t-1}-\bm{G}_{t}\right\|_{2}$ $\leq L\left\|\bm{W}_{t-1}-\bm{W}_{t}\right\|_{2}$ (11)
$=L\left\|\eta\bm{G}_{t-1}^{proj}\right\|_{2}\leq L\eta\left\|\bm{G}_{t-1}\right\|_{2}.$
Consequently, the optimal subspace $\bm{P}_{t}^{\natural}$ is also close enough to the previous optimal subspace $\bm{P}_{t-1}^{\natural}$:
$\left\|\bm{P}_{t-1}^{\natural}-\bm{P}_{t}^{\natural}\right\|_{F}\leq\frac{r\kappa_{1}+\min(m,n)-r}{\kappa_{r}-1}L\eta.$ (12)
When $t\mod(\lambda\times T_{u})=0$, the proposed algorithm updates $\bm{P}_{t}=\bm{P}_{t}^{\natural}$. Otherwise, we show that $\bm{P}_{t}$ highly approximates $\bm{P}_{t}^{\natural}$,
$\left\|\bm{P}_{t}-\bm{P}_{t}^{\natural}\right\|_{F}\leq\left\|\bm{P}_{t-1}^{\natural}-\bm{P}_{t}^{\natural}\right\|_{F}+\left\|\bm{P}_{t}-\bm{P}_{t-1}^{\natural}\right\|_{F}$ (13)
$\leq\left\|\bm{P}_{t-1}^{\natural}-\bm{P}_{t}^{\natural}\right\|_{F}+\left\|\bm{P}_{t}^{\rm MSE}-\bm{P}_{t-1}^{\natural}\right\|_{F}+\left\|\bm{P}_{t}^{\rm SIM}-\bm{P}_{t-1}^{\natural}\right\|_{F}$
$\leq\left\|\bm{P}_{t-1}^{\natural}-\bm{P}_{t}^{\natural}\right\|_{F}+(2-\beta)\left\|\bm{P}_{t-1}-\bm{P}_{t-1}^{\natural}\right\|_{F}$
$\quad+(1-\beta)\sum_{i=2}\beta^{i}\left\|\bm{P}_{t-i}-\bm{P}_{t-1}^{\natural}\right\|_{F}$
$\leq\left\|\bm{P}_{t-1}^{\natural}-\bm{P}_{t}^{\natural}\right\|_{F}+2\left\|\bm{P}_{t-1}-\bm{P}_{t-1}^{\natural}\right\|_{F}$
$\quad+(1-\beta)\sum_{i=2}\beta^{i}\left\|\bm{P}_{t-i}-\bm{P}_{t-1}\right\|_{F}$
$\leq 2^{(t\mod(\lambda\times T_{u}))}\frac{r\kappa_{1}+\min(m,n)-r}{\kappa_{r}-1}L\eta$
1.3 Hyper-parameter Settings.
Table 1 presents the detailed hyper-parameters used in all experiments of COAP. Except for the Update Interval ($T_{u}$) and Re-project Factor ($\lambda$) specific to COAP, all other parameters remain consistent with the corresponding baselines. The optimizer, rank settings, learning rate, GPU configuration, and mixed precision are listed for each experiment. Hyper-parameters for different models, including LDM, SiT-XL/2, ControlNet SDXL, and LLaMA series, are provided to ensure reproducibility.
Table 1: Table 1: Hyper-parameter settings for all experiments.
Optimizer Exprement
Rank
$r$
Rank Ratio
$\alpha$
Update Interval
$T_{u}$
Re-project Factor
$\lambda$
Learning Rate
$\eta$
GPU Mixed-precision
Adafactor (COAP) Table 1. Pre-training LDM - 2 16 10 $2\times 10^{-5}$ 8$\times$×V100 FP32
Table 2. Pre-training SiT-XL/2 512 - 200 5 $1\times 10^{-4}$ 8$\times$×H100 FP16
Table 3. ControlNet SDXL - 2 8 10 $1\times 10^{-5}$ 8$\times$×H100 BF16
- 4 8 10 $1\times 10^{-5}$ 8$\times$×H100 BF16
- 8 8 10 $1\times 10^{-5}$ 8$\times$×H100 BF16
AdamW (COAP) Table 1. Pre-training LDM - 2 16 10 $2\times 10^{-5}$ 8$\times$×V100 FP32
Table 2. Pre-training SiT-XL/2 512 - 30 10 $1\times 10^{-4}$ 8$\times$×H100 FP16
Table 5. Pre-training LLaMA-1B 512 - 40 5 $1\times 10^{-2}$ 8$\times$×H100 BF16
Table 5. Pre-training LLaMA-7B 1024 - 100 1 $5\times 10^{-3}$ 8$\times$×H100 BF16
Table 6. Fine-tuning LLaVA-v1.5-7B - 4 32 1 $2\times 10^{-5}$ 1$\times$×A100 BF16
1.4 Adafactor with COAP.
Adafactor is an adaptive gradient optimization algorithm designed to reduce memory usage while maintaining training efficiency. Unlike Adam, which stores full-matrix second-moment estimates, Adafactor factorizes the second-moment accumulation to save memory, making it particularly suitable for training large-scale models. However, first-moment remains crucial for stabilizing training and accelerating convergence. Without it, optimization can become unstable, especially in deep networks where gradients vary significantly. The first moment helps smooth updates and improves the overall training dynamics. Algorithm 2 describes the Adafactor-based training procedure, by projecting the gradient and its first-moment estimate into a lower-dimensional subspace, we effectively compress the momentum while retaining its benefits. This approach allows us to maintain training stability with reduced memory overhead, making it highly efficient for large-scale models.
2       Betas $[\beta_{1},\beta_{2}]$, Update interval $[\lambda,T_{\rm u}]$, Decay rate $\gamma$.
    $\bm{R}_{0}\in\mathbb{R}^{m\times 1}\leftarrow 0$, $\bm{C}_{0}\in\mathbb{R}^{1\times r}\leftarrow 0$.
4   Randomly Initialize: $\bm{P}_{0}\in\mathbb{R}^{n\times r}$
[…]
6for  $t~{}{\rm in}~{}[1,2,\cdots]$ do
7            Compute: gradient $\bm{G}_{t}$ of $\bm{W}_{t}$ in the loss function.
8      if $t\mod T_{\rm u}=0$ then
9             if $t\mod(\lambda\times T_{\rm u})=0$ then
10                         $\triangleright$ Occasional Low-cost SVD
11                  Compute: $\bm{P}_{t}\leftarrow(\bm{P}_{t-1},\bm{G}_{t})$
12
13            else
14                         Update:$\bm{P}_{t}\leftarrow(\bm{P}_{t-1},\bm{G}_{t},\bm{M}_{t-1})$$\triangleright$ Eqn. 2
15
16
17      else
18             $\bm{P}_{t}\leftarrow\bm{P}_{t-1}$
[…]
20      $\beta_{2}\leftarrow 1-t^{\gamma}$
21      $\bm{G}_{t}^{\rm proj}\leftarrow\bm{G}_{t}\bm{P}_{t}$
23      $\bm{R}_{t}=\beta_{2}\bm{R}_{t-1}+(1-\beta_{2})\cdot\mathrm{Sum}({\bm{G}_{t}^{\rm proj}}^{2},-1)$
24      $\bm{C}_{t}=\beta_{2}\bm{C}_{t-1}+(1-\beta_{2})\cdot\mathrm{Sum}({\bm{G}_{t}^{\rm proj}}^{2},-2)$
25      $\hat{\bm{V}_{t}}=\sqrt{\frac{\mathrm{Mean}(\bm{R}_{t},-1)}{\bm{R}_{t}\bm{C}_{t}}}$
26
[…]
28      $\Delta\bm{W}_{t}^{\rm proj}\leftarrow\beta_{1}\bm{M}_{t}^{\rm proj}+(1-\beta_{1})\eta\hat{\bm{V}_{t}}\odot\bm{G}_{t}^{\rm proj}$
[…]
30      $\bm{W}_{t}\leftarrow\bm{W}_{t-1}-\Delta\bm{W}_{t}^{\rm proj}\bm{P}_{t}^{\top}$
31
32   Return: updated $\bm{W}$
Algorithm 2 Adafactor with COAP
@@ 2 Experimental Results
The COAP optimizer is versatile and can be applied to various model training scenarios. We have provided training results for DDPM and ControlNet, and this method is also suitable for other application scenarios [62, 63, 65, 55].
@@ 2.2 Qualitative Comparisons
We present qualitative results, showcasing images generated by models trained with COAP and other optimizers. These results provide a visual comparison that complements the quantitative analysis in the main paper. Comparisons for DDPM (CIFAR-10), DDPM (CelebA-HQ), LDM, SiT-XL/2, and ControlNet-XL are detailed in Tables 3456, and 7, respectively.